wat is 8x25+50-25/5 =?

Grade:12

8 Answers

Prajwal Kavad
8 years ago
245 is the answer
mamam
8 years ago
thank you
V. Harshit -
8 years ago
245 is the answer
 
Adarsh
8 years ago
 8x25+50-25/5 =200+50-5=245 
Abishek arun
8 years ago
According to the BODMAS rule
8x25+50-25/5 
first divide
8x25+50-25/5
8x25+50-5
secondly multiply
8x25+50-5
200+50-5 = 245
BODMAS rule follows as
brackets then off then divide  then multiply then add then subtract
hope you understand 
please approve my answer
deapan
8 years ago
answer is 245 by following the BODMAS rule
please approve my answer
Hinal Srivastava
8 years ago
 
According to the BODMAS rule
8x25+50-25/5 
first divide
8x25+50-25/5
8x25+50-5
secondly multiply
8x25+50-5
200+50-5 = 245
Hope it helps!!!
SEELA YUVARAJU
8 years ago
245